a person strikes a nail with a hammer of mass 500 g moving with a velocity of 10 m/s . the hammer comes to rest in 0.01 s after striking the nail . calculate the force exerted by hammer on nail. and the distance moved by the nail in to the plank.

Grade:9

1 Answers

Aarti Gupta
9 years ago
Given-
Mass of hammer = 500g or 500/1000 =0.5 Kg
Initial velocity,u = 10m/s
Final velocity,v = 0 m/s as hammer comes to rest
Time,t = 0.01s
Force = ?
Distance,s = ?
Solution-
Using the equation v = u + at
where,a = accelaration
Putting the respective values,
0 = 10 + a*0.01
0.01a = -10
a = -10/0.01 = -1000 m/s2 (minus as retardation)
Now, we know that ,
F = m * a = 0.5 * 1000
So, F = 500N
Now as v2 = u2 + 2as
02 = 102 + 2* -1000 * s
Thus, 2000s = 100
or s = 0.05 m